OpenStudy (anonymous):

Choose the polynomial that is written in standard form. xy3 + 4x5y + 10x3 x4y2 + 4x2y + 8x -9x4y2 + 4x3y5 + 10x2 x6y3 + 4x4y8 + x2

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@jim_thompson5910 No sir. Not a clue.

jimthompson5910 (jim_thompson5910):

to find the degree of something like xy^3, you add the exponents

OpenStudy (anonymous):

So the degree of 2x^4+5xy^2= 6?

jimthompson5910 (jim_thompson5910):

no you add the exponents for the monomial

jimthompson5910 (jim_thompson5910):

degree of 2x^4 is 4 degree of 5xy^2 is 3 (since 1+2 = 3)

jimthompson5910 (jim_thompson5910):

5xy^2 is the same as 5x^1y^2 add the exponents: 1+2 = 3 degree of 5xy^2 is 3

jimthompson5910 (jim_thompson5910):

What's the degree of 4x^5y ?

OpenStudy (anonymous):

7?

OpenStudy (anonymous):

Wait no 6.

jimthompson5910 (jim_thompson5910):

good, 5+1 = 6

jimthompson5910 (jim_thompson5910):

find the degree of each monomial

jimthompson5910 (jim_thompson5910):

the polynomial where you have the degrees decrease is the one written in standard form

OpenStudy (anonymous):

So in this particular case the answer would be D?

OpenStudy (anonymous):

Can you help me with another?

jimthompson5910 (jim_thompson5910):

Notice x^6y^3 has degree: 6+3 = 9 4x^4y^8 has degree: 4+8 = 12 so choice D is not the answer because the degrees are supposed to decrease

OpenStudy (anonymous):

Ohhhh so B?

jimthompson5910 (jim_thompson5910):

much better

OpenStudy (anonymous):

Okay great! Thank you so much!! Can you help me with another one please sir?

jimthompson5910 (jim_thompson5910):

sure, what's your question?

OpenStudy (anonymous):

Simplify: (6x2 - 3 + 5x3) - (4x3 - 2x2 - 16) 2x3 - x2 + 21 -2x3 + x2 - 21 x3 + 8x2 + 13 -x3 - 8x2 + 13

OpenStudy (anonymous):

I put C and was hoping you could verify.

jimthompson5910 (jim_thompson5910):

C is correct

OpenStudy (anonymous):

Great! Thanks!! You dont know how much I appreciate that!!